OpenStudy (anonymous):

Average: 60 Standard diviation: 11.35 What is the probability that a randomly selected monthly cell phone bill is between $45 and $70?

OpenStudy (amistre64):

so.... same concept as last time; but we want the probability from defined points

OpenStudy (amistre64):

z1 = (45 - 60)/11.35 z2 = (70-60)/11.35

OpenStudy (amistre64):

if we use the table i posted in the last one; we would simply then subtract: z2 - z1

OpenStudy (amistre64):

or rather tthe probability values associated with z2 from those of z1

OpenStudy (amistre64):

z1 = -1.32 P(z1) = .4066 z2 = 0.88 P(z2) = .8106

OpenStudy (amistre64):

P(z2) - P(z1) = .404 then if i did it right

OpenStudy (anonymous):

shouldn't it be .3106? for z2

OpenStudy (amistre64):

no, since it is to the right of the mean; z is positive; it is greater than .5 but let me dbl chk to see that I got the right p-value for it

OpenStudy (amistre64):

z=.88 lines up with .8 on the left and .08 from the top to meet at .8106

OpenStudy (amistre64):

on your table it appearts that they measure FROM the mean so you need to add .5 to it

OpenStudy (amistre64):

<---------(0).........(.88) -----------> .5 +.3108

OpenStudy (anonymous):

.404 wasn't the answer, i can't figure it out

OpenStudy (amistre64):

are the zscores good then?

OpenStudy (anonymous):

i think so?

OpenStudy (amistre64):

z2 = .88 ; 10/11.35 z1 = -1.32; -15/11.35 <--------(-1.32).............(0).................(.88)---------> .4066 .5 +.3108

OpenStudy (amistre64):

might be .0934 if i read the wolf right

OpenStudy (amistre64):

.81056-.0934 then

OpenStudy (amistre64):

cannot type today

OpenStudy (amistre64):

.7172 perhaps?

OpenStudy (amistre64):

.9066 - .5 = .4066 from the mean, which if i see it right is good; the other is .3108 from the mean .4066 +.3108 ------- .7174 is the total probability to the left and right of the mean ... i was just intepreting it wrong